OpenStudy (anonymous):

Calculate the number of moles of helium that contains 3.25 x 10^21 atoms.

OpenStudy (anonymous):

Without numbers, can you figure this out. Always start with given value \[\frac{ atoms}{ 1}\times \frac{ mol }{ atoms }\]

OpenStudy (anonymous):

3.25 x 10^ 21 atom x (1mol/6.022 x 10 23 atom)

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

\[3.25 x 10^{21} atom x \frac{ 1 mol }{ 6.022 x 10^{23} atom }\]

OpenStudy (anonymous):

Yes, that is right. Sorry I was helping someone.

OpenStudy (anonymous):

And I get .539687811

OpenStudy (anonymous):

Units would be in moles of helium.

OpenStudy (anonymous):

So .539687811 mol He?

OpenStudy (anonymous):

It has to be 3 sf since the given value only has 3. Remember that your answer has a 0 in the front. It does not count as a sf.

OpenStudy (anonymous):

.540 mol He?

OpenStudy (anonymous):

Yes Nice rounding!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

It's saying the answer is wrong. I did it right...Right?

OpenStudy (anonymous):

I did not do the calculations since you got the right conversion. You did the calculation. Do the calculation again.

OpenStudy (anonymous):

I think you put 3.25*10^23 instead of 3.25*10^21

OpenStudy (anonymous):

I did it again and got the same answer.

OpenStudy (anonymous):

Write down what you calculate then.

OpenStudy (anonymous):

\[3.25\times10^{21} atomtimes \frac{ 1mol }{ 6.022 \times10^{23 } }atom\]

OpenStudy (anonymous):

3.25 x 10^21 atom x

OpenStudy (anonymous):

Dont forget parentheses. (3.25*10^21)/(6.022*10^23)

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Got .00540 mol He

OpenStudy (anonymous):

So what did you do wrong earlier?

OpenStudy (anonymous):

Forgot parentheses or wrong exponent?

OpenStudy (anonymous):

I put it in the calculator without parentheses

OpenStudy (anonymous):

Another way is multiply 3.25*10^21 and lets called that x then multiply 6.023*10^23 and call that y. Then just do x/y if you arent comfortable with parentheses.